How long will it take $3000 to double if it is invested in an account that pays 3% compounded continuously?
Answer by Alan3354(69443)

FV = PV*e^(rt)
2PV = PV*e^(0.03t)
e^(0.03t) = 2
0.03t = ln(2)
t = ln(2)/0.03
t =~ 23.1 years